MOBY-L279 experiment with MOBYrefresh - BS03 & RS04 -Pre-MOBY267 deployment with BS/RS- NEW PROCESSING

BS03 postcal data collected after deployment on MOBY267 with the RS04.

Art had the new h5 file being created to save the auxilliary data during Mikes data collection. Images are still saved as FITS files.

BLUE jumper cables RED jumper cables
Track Sensors filter     Track Sensors filter    
2 Es 0.5ND     2 Es 0.5ND & 1.0ND    
6 LuTop none     6 LuTop none    
9 LuMid none     9 LuMid none    
13 LuBot none     13 LuBot none    

Problems seen (and solved) and good things:

  1. RS only: For lower pixels the net values are negative. When looking at the dark and light values on the track that is illuminated have the lights lower than the darks, only for the track illuminated. The problem is seen in all the tracks but not all the time.

Log Sheets

Photos

MOBY265 postcal Es\Lu rsp BS & RS

  BS RS
Wavelength cals P(2,:) = [-4.6267e-09 1.872036e-05 0.3363279 337.6345];
P(6,:) = [-4.422466e-09 1.827874e-05 0.3366641 337.5965];
P(13,:) = [-3.387284e-09 1.677827e-05 0.3371351 337.7383];
P(2,:) = [-5.389367e-09 2.018959e-05 0.3856049 502.5337];
P(6,:) = [-3.157798e-09 1.644415e-05 0.387208 502.4158];
P(13,:) = [-5.427442e-09 2.018914e-05 0.3856662 502.4122];
int time cal (from L274) Equation: a+b*x+c*x*log(x)+d./x^(0.5)
a,b,c,d = [0.04654 0.94308 0.0029037 0.0052378];
Equation: a+b*x+c*x*log(x)+d./x^(0.5)
a,b,c,d = [0.036755 0.94299 0.0026703 0.0088647];
Track Definitions
Track#: [ 1   2   3   4   5   6   7   8   9  10  11  12  13  14]
 Left:  [ 1  80 153 229 297 367 447 516 581 661 728 800 873 950]
Right:  [79 152 228 296 366 446 515 580 660 727 799 872 949 1024]
Track#: [ 1   2   3   4   5   6   7   8   9  10  11  12  13  14]
Left:   [ 1  74 150 223 296 370 441 512 585 658 731 802 875 953]
Right:  [73 149 222 295 369 440 511 584 657 730 801 875 952 1024]
All the raw data files and their images , Table of KEYWORDS variable
Page Number
Link
Description
Date
1.01
Day 01 - 29 Mar Raw Data: BS & RS | Table : BS & RS
DAY01 - led testing of new software and BS trk12 shutter and RS trk 4 shutter (BS cfg 007, RS cfg 004)
30 Mar 2019
1.02
Day 02 - 30 Mar Raw Data: BS & RS | Table : BS & RS
DAY02 - led test of Es (BS cfg 007, RS cfg 004)
30 Mar 2019
1.03
Day 03 - 1 Apr Raw Data: BS & RS | Table : BS & RS
DAY03 - System response marathon (Es and all three Lus) (BS cfg 008, RS cfg 005)
1 Apr 2019
1.04
Day 04 - 2 Ap rRaw Data: BS & RS | Table : BS & RS
DAY04 - After ND addition, 2nd Es cal and some wavecal on Trk 2, 6 and 13 (BS cfg 009, RS cfg 006 - final for MOBY267)
2 Apr 2019
Daily data sets
All Days
2.01 Track movement: BS & RS Checking Track movement  
Day 4 data
5.01 Es wavlength cal BS & RS Es wavecals (HgA,Ne) 2 Apr 2019
5.02 LuTop wavlength cal BS & RS LuTop wavecals (Ar,HgA,Ne,Kr,Xe) 2 Apr 2019
5.03 LuBot wavlength cal BS & RS LuBot wavecals (HgA,Ne) 2 Apr 2019
5.04 comp wave cals BS & RS Compare the L274 and L279 wavecals  
5.05 Es ND filter installation BS & RS Comparing the Es data before and after installtion of ND filters 2 Apr 2019
5.06 Es/Lu system response BS & RS Es/Lu system response data from day 3 (Lu) and 4 (Es) 2 Apr 2019
5.07 Estimate int times -> BS & RS Estimate the in-water signals at 10 seconds int time 16 Apr 2019
5.08 balancing test -> BS & RS Make a try and picking ND filters for balancing Es and Lu 17 Apr 2019
automated graphs of data over many days
20.01 UV nets-> BS & RS RS low pix data are negative- We found some nevative values in the low pixel on H19-01 and are checking we see them here 31 Jan 2019
20.02 darks-> BS link & RS link Darks from all experments 17 Apr 2019
DAY 1 EMAIL from Mike - LED data

Hi Steph,
There are some files from last night for you at /ftp1/Mike/L279 !
The log sheets are not complete - I need to figure out cfg#s for the BS03 & RS04,
so those will get updated sometime soon.
These scans just tested the shutters with bogus return states,
and acquired the first HDF file with BS & RS aux data.
Luck!
MF

DAY 2 EMAIL from Mike - LED more checks
Hi STePh,
There are some more data for you a the NAS /ftp1/Mike/L279/ !

Pre-M267 / pre-L279 Day02 was only the Es FO connected @ Trk02.
I took 1x scan set of LED Flashlight input.
There are some photos, showing the 2x tight coils in the FO
near the FO block at the Splitter housing.

Art said he updated the asg_interactive software to fix an
error you found in the aux file.

I also re-scanned and re-named the log sheets from Day01.

Tomorrow we'll try to get the 3x Lu FOs connected.
Aloha, MF

DAY 3 EMAIL from Mike - 1 Es and 1 Lu
Hi Steph,
Thanks for your work on days 1:2 data!
I have not had a chance to look at them, but from your emails
it sounds like things are kinda making sense (?).

There are day3 data at the NAS now!
There were 2x log sheets - 1 for Es cal, 1 for Lu cals.
I also moved lamp-monitor files to \L279\bs3cal\ = rb67BSRS*.txt

If possible, could you make the raw & net & rsp .mat files for me
like you did for the post-M265 BS & RS cals?
From these I think I can guess at what attenuation needs to be
made to the Es responses - i.e. this will help determine what
Neutral Density filters to put on the Es tracks.

It looks like we have to load the buoy on the torpedo boat tomorrow.
So everything will have to be finished today.  Which means I doubt
there will be time for a full set of wavelength & exp time cals.
We'll see how that goes...
Cheers, MF
DAY 4 EMAIL from Mike - Es ND filters added and Es rsp and wavecals
hI StePh,
Day4 data are at the NAS.  Those were Es cal #2 & wave cals.
New cfgs: BS03c009 Es got a 0.5 ND filter, RS04c006 Es got a 1.5 ND filter.
The 1.5 ND filter was actually a combination of a 0.5 & 1.0 ND filters.
I have some more photos that are not yet ready to give to you.
I.E. photos of the ND filters, FO pucks, OM mounting, etc.

I found the missing BS putty log, so those log files should all be OK.

As we "speak" the final button-up of MOBY267 should be in progress.
The plan is to load the boat today (Tue), depart tonight, be at MOBY
tomorrow morning & tow the ole M266 to Kaumalapau harbour.
Thursday we use the crane at the dock to swap the buoys,
then tow the new M267 back out the mooring.

Yikes, I almost forgot aboot the MOS response!  There's are new
byy_rsp7.m/at files at NAS /ftp1/Mike/cal_hist/.
I don't think my .xlsx config files are up to date though...
Aroha, MF
ND filter emails after day 4 data collection

Hi Steph,
I thought you might like to have the Transmission data from Thorlabs
for the 2x ND filters which got added to the BS,RS Es sensors:
<attached: NDUV*.*>
MF

NDUV_OD05.txt NDUV_OD1.txt
NDUV_OD05.gif NDUV_OD1.gif
NDUV_OD05.xlsx NDUV_OD1.xlsx

ND filter emails after day 4 data collection

Hi Steph,
Here is more "documentation" of BS,RS Es ND filters:
<attached: 20190401_M267-BS03cfg009-Es-ND-filter_1.jpg>
<attached: 20190401_M267-RS04cfg006-Es-ND-filter_1,  2.jpg>
MF

20190401_M267-BS03cfg009-Es-ND-filter_1.jpg
20190401_M267-RS04cfg006-Es-ND-filter_2.jpg
20190401_M267-RS04cfg006-Es-ND-filter_1.jpg

Mikes email on pre-M267 BSRS Es w/ ND filters

Hi Steph,
I've been wanting to make this plot:
<attached: preM267_BSRS_Es_resp_1.png>

Then I poked around in your www L279 plots and found that you'd already done just that!

I'm very happy that this worked out the way it did.
Here are my "Expected" in-water 10 sec exposure time signals:
<attached: preM267_BSRS_Expected_12.png>

Now we just need to get some in-water data...
MF

es1 est1 est2